20201128 LBTB Science

00:18 Open
  • Sky flats first
  • To dark for flats

00:31 Pointing and collimation

00:50 OCSO
  • One set with NCPA off
  • GOPT is not measuring a ratio
  • Optical Diag. for Gopt log says 1000 frames are being saved
  • Find in /tmp/od_gopt.stdout
FITSIO status = 105: couldn't create the named file
failed to create new file (already exists?):
/local/aomeas/adsec_data/20201129/Data_20201129_250000/Timestamp_20201129_250000
.fits
  • Should make the gopt routine do a rm -f so it does not ask us to delete file
  • I did a rm -fr /local/aomeas/adsec_data/20201129/Data_20201129_250000/
  • OSCO_Modal.jpg: I can see the mode applied
    OCSO_Modal.jpg

01:23 PresetAO failed
  • BCU47 and Tip Tilt power off/on on Power controller HW GUI
WfsArbitrator_94205|ERR|       496|2020-11-29 01:23:06.075076|             MAIN > Cannot turn off Bcu47 RESET-Active Low. Reason: Timeout waiting for powerboard.L.RESET_BCU_47.CUR - TIMEOUT_ERROR (-1000) GENERIC_SYS_ERROR
  • Power_screen.jpg: Note, this is after the power off/on of box47. Before just Bcu47 and Tip Tilt power was off and all else on.
    Power_screen.jpg
  • When to set binning from HW GUI
    • Setting binning: 16
    • Error: Timeout waiting for techviewctrl.L.STATUS - TIMEOUT_ERROR=
  • Tried turning box 47 power off/on
  • Powered off, w_stop_all, w_start_all
  • First time I powered off I received error
  • Power_off_error.jpg:
    Power_off_error.jpg
  • image
  • Operate
  • Everything starts up fine.
  • Looked like Bcu47 powered on with Aux Unit Power on
    • Next time I'll try Aux Unit power to get Bcu47 to power on
  • BaysideX did not home. Rotator not at park. Josh move rotator and X homed
    • Need to add this to IT list

02:00 All Running now
  • Preset
  • Camera lens at -100, Pupil will not converge
  • Rotator at 160
  • Jochen change script PA by +90 and resend
  • TN 20201129_021006

02:12 New preset
  • pupil aligned with x = -85.5, y = -37.0
  • Loop closed with no problems....other than...
  • GOPT is not working
    • Still see "file exists" error in /tmp/od_gopt.stdout
    • Even if I remove this directory, it writes once and then the error occurs and timestamps of files do not change
  • Bin 3 so running without Gopt and NCPA
  • TN 023129, 023525
  • bin3_modal.jpg:
    bin3_modal.jpg

03:10 Try to get GOPT to work
  • Invert matrix via elab
$ idl
IDL> ao_init
AO> a = getaoelab('20201129_023129')
AO> help, (a->control())->c2m()
AO> exit
  • GOPT now works.
    • Not getting "file exist" error in od_gopt.stdout
    • ratio is now 0.02-0.03. Seems really low
    • can not see modulation signal in modalplot
    • TN 031910, 032129
  • from May 18, 2016
Hi Doug,
from the logs it appears like a known problem: when the gopt first encounters a new modal basis (KL_v16) in this case, the elab_lib needs to invert the m2c matrix to obtain a c2m, but for some reason the IDL inversion routine crashes when called from gopt.
The solution is to open a TN with the elab_lib and make it calculate the c2m, for example asking for (a->control())->c2m()

I have just done that with a fresh TN, so try again the gopt and see if it repeats

Alfio

03:22 Power outage at LBTO.
  • Click AMBIENT on WFS OCAM
  • set Ocam EM gain to 1
  • ADSEC in failure
  • wind low so will restart here ~1.0 m/s
  • HBS off so can't move anyway
  • Click off OCAM Chiller to AMBIENT until we get everything back up
  • AdSec is now restarted and set. No Problems
  • Current values on AdSec Mirror GUI are all grey

04:58 Power back up and now colimated
  • OK for center star, pupil and flux
  • noise stripe down the middle of the OCAM
  • set binning to 4 then back to 3. Looks good now.
  • Optimize gain ok
  • GOPT 0.9
  • Bin 3, 918 Hz
  • NCPA On
  • TN 051752, 051812, 052216, 054145, 055315
  • 14.9 mag
  • bin3_gopt.jpg:
    bin3_gopt.jpg
  • bin3_mode_gopt_on.jpg:
    bin3_mode_gopt_on.jpg

06:32 AO Loop open due to skipframe
  • reclose loop, GOPT on, NCPA on
  • TN 065448

07:00 Done, hand over to PEPSI

-- DougMiller - 29 Nov 2020

I Attachment Action Size Date Who Comment
OCSO_Modal.jpgjpg OCSO_Modal.jpg manage 62 K 29 Nov 2020 - 02:46 DougMiller  
Power_off_error.jpgjpg Power_off_error.jpg manage 24 K 29 Nov 2020 - 02:46 DougMiller  
Power_screen.jpgjpg Power_screen.jpg manage 133 K 29 Nov 2020 - 02:46 DougMiller  
bin3_gopt.jpgjpg bin3_gopt.jpg manage 61 K 29 Nov 2020 - 05:25 DougMiller  
bin3_modal.jpgjpg bin3_modal.jpg manage 60 K 29 Nov 2020 - 02:47 DougMiller  
bin3_mode_gopt_on.jpgjpg bin3_mode_gopt_on.jpg manage 57 K 29 Nov 2020 - 05:25 DougMiller  
Topic revision: r3 - 29 Nov 2020, DougMiller
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback